OpenStudy (anonymous):

A spherical balloon is inflated with a gas at a rate of 240 cubic centimeters per minute. How fast is the radius of the balloon increasing at the instant the radius is 30 cm?

OpenStudy (jtvatsim):

Welcome to OpenStudy! Is this a question from Calculus class?

OpenStudy (anonymous):

yes it is.

OpenStudy (freckles):

so we have a sphere do you know the formula for a sphere ?

OpenStudy (freckles):

well the volume of a sphere that is

OpenStudy (anonymous):

Volume- (4/3) (pi)(r)^3

OpenStudy (freckles):

ok differentiate that equation w.r.t r

OpenStudy (freckles):

\[V(r)=\frac{4}{3}\pi r^3 \] that is find: \[\frac{dV}{dr}\]

OpenStudy (freckles):

do you know how to find: \[\frac{d(r^3)}{dr}=?\]

OpenStudy (anonymous):

not really

OpenStudy (freckles):

have you guys talked about power rule ?

OpenStudy (anonymous):

no. my teacher does not explain anything right

OpenStudy (freckles):

have you guys talked about definition of derivative?

OpenStudy (freckles):

could you evaluate this limit: \[\frac{d(r^3)}{dr}=\lim_{h \rightarrow 0} \frac{(r+h)^3-r^3}{h}\]

OpenStudy (freckles):

try expanding (r+h)^3 first